OpenStudy (destinykiara99):

What are the solutions to the following system of equations? Select the correct answer below. y = x2 − 2x + 4 8x + y = 20

OpenStudy (kinged):

Make the second equation in terms of y.

OpenStudy (kinged):

Then equate both equations and then solve for x.

OpenStudy (destinykiara99):

y=x^2-2x+4 y=20-8x x^2-2x+4=20-8x

OpenStudy (destinykiara99):

Solve for X?

OpenStudy (destinykiara99):

@kinged

OpenStudy (will.h):

you still need help?

OpenStudy (destinykiara99):

Yes I do

OpenStudy (destinykiara99):

I know the X values but not the Y so I have it narrowed down to two possible answers

OpenStudy (will.h):

ok, 1st of all there are ways to solve a system of equations. let's first correct the shape of the equations. \[y=x^2 -2x + 4\] \[y=-8x + 20\] Good so far?

OpenStudy (destinykiara99):

Yes, I already have that. I know that when you make the two equations equal to each other X=2 and X=-8

OpenStudy (will.h):

can you solve those equations using substitution method? Like instead of y in the second equation we can plug x^2 -2x + 4 Like this \[x^2 - 2x + 4 = -8x + 20\]

OpenStudy (will.h):

AND YES I JUST SOLVED IT YOUR ANSWER IS CORRECT YOU CAN PLUG X IN ANY EQUATION TO FIND Y HOPE THAT HELPS

OpenStudy (destinykiara99):

Thank you so much, I got it now :)